let g(S) = square root of s(4-S^2) find g'(s)
Answer by stanbon(75887)

let g(S) = square root of s(4-S^2) find z'(s)
-----
g(S) = S^(1/2)*(4-S^2)
------
g'(S) = (1/2)S^(-1/2)*(4-S^2) + (-2S)*S^(1/2)
-----------------
g'(S) = (4-S^2)/[2S^(1/2))] - 2S^(3/2)
-------------------------
Cheers,
Stan H.
----------------